OpenStudy (gamenerd123):

The sum of three consecutive even integers is 36. Let x represent the first integer, x + 2 represent the second integer, and x + 4 represent the third integer. Which equation can be used to find the first integer? 3x = 2x + 6 2x + 6 = 36 3x + 36 = 2x + 6 3x + 6 = 36

OpenStudy (gamenerd123):

Help please?

OpenStudy (mhchen):

"Let x be the first integer" So x is the first integer. "Sum...is 36" First + Second + Third = 36 (x) + (x+2) + (x+4) = 36 x + x + 2 + x + 4 = 36 <-- Can you simplify this

OpenStudy (gamenerd123):

Not on my own

OpenStudy (gamenerd123):

I'm not very familiar with simplifying integers

OpenStudy (gamenerd123):

\[2x + 6 = 36?\]

OpenStudy (mhchen):

So treat 'x' as a pencil or somethin. If you have a pencil + pencil + pencil, that's 3 pencils. So x + x + x = 3x Makes sense? So x + x + 2 + x + 4 = 3x + 2 + 4 Right? And you know how to do 2 + 6 ofc 3x + 2 + 4 = 3x + 6 And from our original equation, the left side is 36. So if we combine them: 3x + 6 = 36.

OpenStudy (triciaal):

simplifying integers just mean add the numbers

OpenStudy (mhchen):

*2 + 4 I mean, not 2+6 srry

OpenStudy (triciaal):

add like terms

OpenStudy (gamenerd123):

\[3x+6=36?\]

OpenStudy (gamenerd123):

I'm sorry for giving a hard time

OpenStudy (triciaal):

you are not giving a hard time! you are here to learn and we are here to help be positive

OpenStudy (gamenerd123):

So what I posted was the right answer?

OpenStudy (triciaal):

yes

OpenStudy (gamenerd123):

*plz be right plz be right plz be right* YAAAY!!!!

OpenStudy (triciaal):

important to understand steps not just get a correct answer

OpenStudy (gamenerd123):

Can you help me with one more please?

OpenStudy (gamenerd123):

Oh, so far I comprehend it, I also took notes

OpenStudy (triciaal):

when you understand the steps you will get the correct answer except for simple mistakes in calculations

OpenStudy (gamenerd123):

Oh

OpenStudy (gamenerd123):

I'll remember that

OpenStudy (triciaal):

open a new post with the new question

OpenStudy (gamenerd123):

Ok